OpenStudy (anonymous):

Help!! I will medal and fan :) Julia pays a flat rate of $106 for her cell phone and is charged $0.12 for every text she sends. Julia spends at least $142 on her phone bill each month. Which inequality will help her determine the minimum number of texts she sends per month? $106 + $0.12x ≤ $142 $106 + $0.12x ≥ $142 $106 + $0.12x < $142 $106 + $0.12x > $142

jimthompson5910 (jim_thompson5910):

The cost equation is c = 0.12x + 106 which can be rearranged to c = 106 + 0.12x

jimthompson5910 (jim_thompson5910):

where x = # of text Julia sends

jimthompson5910 (jim_thompson5910):

texts*

jimthompson5910 (jim_thompson5910):

so if "Julia spends at least $142 on her phone bill each month.", then we know the cost is at least $142

jimthompson5910 (jim_thompson5910):

so this means \[\Large c \ge 142\] plug in c = 106 + 0.12x to get \[\Large 106 + 0.12x \ge 142\]

OpenStudy (anonymous):

Oh that makes perfect sense. Thank you soooooo much!

jimthompson5910 (jim_thompson5910):

you're welcome